40ml of amixture of hydrogen,CH4 and N2 was exploded with 10ml of oxygen.on cooling, the gases occupied 36.5ml.after treatment with KOH, the volume reduced by 3ml and again on treatment with alkaline pyrogallol,the volume further decreased by 1.5ml.determine the composition of the mixture.

Grade:upto college level

1 Answers

Sunil Kumar FP
9 years ago

KOH solution CO2& SO2
Alkaline Pyrogallol O2

absorbed gas
equation
CH4+2O2 ----CO2 + 2H2O
N2 +O2 -----2NO

volume of CO2=3ml
volume of O2 remaining=1.5ml
volume of NO=36.5-(4.5)=32ml
apply poac volume of N2=16ml
volume of CH4=24ml
percentage composition=16/40*100=40%
CH4=60%